OpenStudy (anonymous):

Solve for the indicated domain: cosx + sqrt3 sinx=1 {0, 2pi}

OpenStudy (anonymous):

write as a single function first

OpenStudy (anonymous):

\[\cos x+\sqrt{3} \sin x=1 ; x \in [0, 2\Pi]\]

OpenStudy (anonymous):

first write \[\cos(x)+\sqrt3\sin(x)\] as a single sine function

OpenStudy (anonymous):

do you know what i mean here?

OpenStudy (anonymous):

Not really.......

OpenStudy (anonymous):

ok you can always turn \[b\cos(x)+a\sin(x)\] in to a single sine function

OpenStudy (anonymous):

we can walk through it if you like you are going to take \[\cos(x)+\sqrt{3}\sin(x)=k\sin(x+\theta)\] what you need is \(k\) and \(\theta\)

OpenStudy (anonymous):

it is a direct consequence of the addition angle formula for sine, but you don't need the derivation, just the mechanics

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[k=\sqrt{a^2+b^2}=\sqrt{1^2+\sqrt{3}^2}=\sqrt{1+3}=\sqrt{4}=2\]

OpenStudy (anonymous):

I got that! Yay

OpenStudy (anonymous):

k now for \(\theta\) use \[\sin(\theta)=\frac{b}{k}\] and here it is important to remember that \(b\) is the coefficient of the COSINE term, so we know \[\sin(\theta)=\frac{1}{2}\]

OpenStudy (anonymous):

also \[\cos(\theta)=\frac{a}{k}=\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

so what it \(\theta\) ?

OpenStudy (anonymous):

.............That's where i got stuck

OpenStudy (anonymous):

you know an angle where the sine is one half and the cosine is root three over two??

OpenStudy (anonymous):

\[\Pi/6\] or 30 degrees

OpenStudy (anonymous):

yeah lets be adults and stick with number rather than degrees so we get \[\cos(x)+\sqrt3\sin(x)=2\sin(x+\frac{\pi}{6})\]

OpenStudy (anonymous):

you can check this is right by using the addition angle formula for the right hand side and see that you get the left hand side if you like now your job is to solve \[2\sin(x+\frac{\pi}{6})=1\] which should take only two or three steps

OpenStudy (anonymous):

you good from there? or still need help?

OpenStudy (anonymous):

oh i was doing something. Back now though @~@

OpenStudy (anonymous):

lol pretty sophisticated problem for high school math especially in the summer

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Question: Do you have to divide both sides by 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So then I get\[\sin x \cos \Pi/6+\sin \Pi/6\cos x\]

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

\[2\sin(x+\frac{\pi}{6})=1\] divide by 2 and get \[\sin(x+\frac{\pi}{6})=\frac{1}{2}\]

OpenStudy (anonymous):

I did, then I expanded the left side

OpenStudy (anonymous):

don't do that that is working backwards

OpenStudy (anonymous):

if \[\sin(x+\frac{\pi}{6})=\frac{1}{2}\] then \[x+\frac{\pi}{6}=\frac{\pi}{6}\] and so \(x=0\) one solution

OpenStudy (anonymous):

Oh. Derp

OpenStudy (anonymous):

then there is probably another for example \(\sin(\frac{5\pi}{6})=\frac{1}{2}\) and so you can solve \[x+\frac{\pi}{6}=\frac{5\pi}{6}\] for \(x)\)

OpenStudy (anonymous):

x= 2pi/ 3 right

OpenStudy (anonymous):

yeah i think that looks good

OpenStudy (anonymous):

Ok thank you so much! I have a test today and I needed help OMG THANK YOU

OpenStudy (anonymous):

yw good luck!